OpenStudy (anonymous):

Simplify -(-y)^2(y)^4 |(-7)(-m)^3(-m^2( (-b^3)(-b^4)(-b^5) (1/8)(-8x) (-9/10x)(20/3x) (-3/4)(-k)(-16k)

jimthompson5910 (jim_thompson5910):

You've posted quite a lot here Did you get anywhere with any of them?

OpenStudy (debbieg):

PEMDAS. Eg. on the first one: \(\large -(-y)^2(y)^4\) Start with the exponents (the only one that you really need to deal with is the\(\large (-y)^2\) ), then multiply left to right. Keep in mind: \(-a^2\) is really just a factor of -1, multiplied by \(a^2\), so if you have something like: \(-a^2a^5\) that is really equivalent to: \((-1)\cdot a^2\cdot a^5=-a^7\)

OpenStudy (zpupster):

hey abby can you try the second one and we will help you

OpenStudy (debbieg):

Or maybe the first one, first. :)

OpenStudy (zpupster):

oh yeah the first one -- at a glance i thought debbie solved it -- sorry debbie